# Higher-Order Regularization Learning on Hypergraphs

### v1

@@ 4.2.1 Convergence of truncated energies in the single Laplacian case
First, we tackle the $\liminf$-inequality. We assume that $(\nu,v)\in\mathrm{TL}^{2}(\Omega)$ and that $(\nu_{n},v_{n})\to(\nu,v)$ in $\mathrm{TL}^{2}$. If $\liminf_{n\to\infty}(\mathcal{S}\mathcal{J})_{n,\{\varepsilon_{n}\},K_{n}}^{(1,\{s\})}((\nu_{n},v_{n}))=+\infty$, then the inequality is trivial. Hence, without loss of generality, let us assume that $\sup_{n\in\mathbb{N}}(\mathcal{S}\mathcal{J})_{n,\{\varepsilon_{n}\},K_{n}}^{(1,\{s\})}((\nu_{n},v_{n}))\leq C$. In particular, this implies that $\nu_{n}=\mu_{n}$, $\langle\phi_{n,\varepsilon_{n},k},v_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}=0$ for all $k>K_{n}$ and $\sup_{n\in\mathbb{N}}\sum_{k=1}^{K_{n}}a_{n,\varepsilon_{n},k}^{s}\langle\phi_{n,\varepsilon_{n},k},v\rangle_{\mathrm{L}^{2}(\mu_{n})}^{2}\leq C$. Since we have $\mu_{n}\to\nu$ weakly (by the $\mathrm{TL}^{2}$-convergence assumption—see Proposition 2.2) and $\mu_{n}\to\mu$ weakly (convergence of the empirical measures), we conclude (by the uniqueness of weak limits) that $\nu=\mu$. We then proceed as in [22, Theorem 2.2].
Let us start by assuming that $\sum_{k=1}^{\infty}\beta_{k}^{s}\langle v,\psi_{k}\rangle_{\mathrm{L}^{2}(\Omega)}^{2}<\infty$. In particular, since $\phi_{n,\varepsilon_{n},k}\to\psi_{k}$ and $v_{n}\to v$ in $\mathrm{TL}^{2}(\Omega)$ [22], we have that $\langle\phi_{n,\varepsilon_{n},k},v_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}\to\langle v,\psi_{k}\rangle_{\mathrm{L}^{2}(\mu)}$ by [36, Proposition 2.6]. Furthermore, by [36, Theorem 1.2], we have $a_{n,\varepsilon_{n},k}\to\beta_{k}$. Now, let $\delta>0$ and pick $K$ such that
@@
For the $\limsup$-inequality, we let $(\nu,v)\in\mathrm{TL}^{2}(\Omega)$. If $(\mathcal{S}\mathcal{J})_{n,\{\varepsilon_{n}\},K_{n}}^{(1,\{s\})}((\nu,v))=\infty$, the inequality is trivial, so we assume that $\nu=\mu$ and $\sum_{k=1}^{\infty}\beta_{k}^{s}\langle\psi_{k},v\rangle_{\mathrm{L}^{2}(\mu)}^{2}<\infty$ or equivalently $v\in\mathrm{W}^{s,2}(\Omega)$ [22]. If we can prove the $\limsup$-inequality on a dense subset of $\{\mu\}\times\mathrm{W}^{s,2}(\Omega)$, namely $\{\mu\}\times\mathrm{C}_{c}^{\infty}(\Omega)$, we can conclude due to [35, Remark 2.7].
@@
where we used the fact that the eigenvalues are ordered for (46) and [87, Lemma 4.19] for (47). We know from [36, Theorem 1.4] that $T_{1}\to 0$ and from the proof of [22, Theorem 2.2] that $a_{n,\varepsilon_{n},K_{n}+1}^{2}\to\infty$ which allows us to conclude that $(\mu_{n},\bar{v}_{n})\to(\mu,v)$ in $\mathrm{TL}^{2}(\Omega)$.
@@
where used [87, Lemma 4.15] for (48), [87, Proposition 4.21] and the fact that $T_{2}\to 0$ for (49). Squaring the last inequality, we obtain the $\limsup$-inequality.
@@
Let us now consider a sequence $(\mu_{n},v_{n})$ minimizing $(\mathcal{S}\mathcal{J})_{n,\{\varepsilon_{n}\},\Psi,K_{n}}^{(1,\{s\})}$ with $\sup_{n\in\mathbb{N}}\|v_{n}\|_{\mathrm{L}^{2}(\mu_{n})}\leq C$. In particular, we note that $\langle\phi_{n,\varepsilon_{n},k},v_{n}\rangle_{n}=0$ for all $k>K_{n}$ and recall that $K_{n}\to\infty$. Therefore, we can apply the same proof as in [22, Theorem 2.2] to show that there exists a converging subsequence in $\mathrm{TL}^{2}(\Omega)$.
Specifically, $\sup_{n\in\mathbb{N}}\|v_{n}\|_{\mathrm{L}^{2}(\mu_{n})}\leq C$ implies that $\sup_{n\in\mathbb{N}}\sum_{k=1}^{K_{n}}\langle v_{n},\phi_{n,\varepsilon_{n},k}\rangle_{\mathrm{L}^{2}(\mu_{n})}^{2}\leq C.$ Hence, by a diagonal procedure, we can find a sequence $n_{m}\to\infty$ such that for every $k$, $\langle v_{n_{m}},\psi_{n_{m},\varepsilon_{n_{m}},k}\rangle_{\mathrm{L}^{2}(\mu_{n_{m}})}$ converges to some coefficient $\gamma_{k}$. By Fatou’s lemma, $\sum_{k=1}^{\infty}|\gamma_{k}|^{2}\leq\liminf_{m\to\infty}\sum_{k=1}^{n_{m}}|\langle v_{n_{m}},\phi_{n_{m},\varepsilon_{n_{m}},k}\rangle_{\mathrm{L}^{2}(\mu_{n_{m}})}|^{2}\leq C$, so we can define $v=\sum_{k=1}^{\infty}\gamma_{k}\psi_{k}\in\mathrm{L}^{2}(\mu)$. Using [22, Lemma 7.7], we obtain a sequence $R_{n_{m}}\to\infty$ such that $\sum_{k=1}^{R_{n_{m}}}\langle v_{n_{m}},\phi_{n_{m},\varepsilon_{n_{m}},k}\rangle_{\mathrm{L}^{2}(\mu_{n_{m}})}\phi_{n_{m},\varepsilon_{n_{m}},k}\to v$ in $\mathrm{TL}^{2}(\Omega)$. We note that $R_{n_{m}}$ can always be picked such that $R_{n_{m}}\leq K_{n_{m}}$. Indeed, the $\mathrm{TL}^{2}$-convergence resulting from [22, Lemma 7.7] holds for any sequence converging to $\infty$ and majorized by $R_{n_{m}}$: therefore, we can always pick $\tilde{R}_{n_{m}}=\min\{R_{n_{m}},K_{n_{m}}\}$.
@@
$+\|\sum_{k=1}^{R_{n_{m}}}\langle v_{n_{m}},\phi_{n_{m},\varepsilon_{n_{m}},k}\rangle_{\mathrm{L}^{2}(\mu_{n_{m}})}\phi_{n_{m},\varepsilon_{n_{m}},k}\circ T_{n_{m}}-v\|_{\mathrm{L}^{2}(\mu)}$
@@
$+\|\sum_{k=1}^{R_{n_{m}}}\langle v_{n_{m}},\phi_{n_{m},\varepsilon_{n_{m}},k}\rangle_{\mathrm{L}^{2}(\mu_{n_{m}})}\phi_{n_{m},\varepsilon_{n_{m}},k}\circ T_{n_{m}}-v\|_{\mathrm{L}^{2}(\mu)}$
@@
where the last inequality follows from the fact that $\sup_{n\in\mathbb{N}}\sum_{k=1}^{K_{n}}a_{n,\varepsilon_{n},k}^{s}\langle v_{n},\phi_{n,\varepsilon_{n},k}\rangle_{\mathrm{L}^{2}(\mu_{n})}^{2}\leq C$ since $(\mu_{n},v_{n})$ are minimizers of $(\mathcal{S}\mathcal{J})_{n,\{\varepsilon_{n}\},\Psi,K_{n}}$ (see also [87, Lemma 4.25]). In order to conclude that $v_{n_{m}}$ to $v$ in $\mathrm{TL}^{2}(\Omega)$, we note the following two facts: the first term in the last inequality tends to 0 as argued in [22, Theorem 2]; the second term tends to 0 since $\sum_{k=1}^{R_{n_{m}}}\langle v_{n_{m}},\psi_{n_{m},k}\rangle_{\mathrm{L}^{2}(\mu_{n_{m}})}\psi_{n_{m},k}\to v$ in $\mathrm{TL}^{2}(\Omega)$. By Proposition 2.6, we know that the limiting point $v$ is a minimizer of $(\mathcal{S}\mathcal{J})_{\infty,\Psi}^{(1,\{s\})}$. ∎
@@ 4.2.2 Spectral convergence of $\mathcal{L}_{n}^{(q)}$
$\mathcal{L}^{(q)}\psi_{i}=\sum_{k=1}^{\lambda_{k}}\Delta_{\rho}^{p_{k}}\psi_{i}=\sum_{k=1}^{q}\lambda_{k}\beta_{i}^{p_{k}}\psi_{i}.$
@@
Again, as $\{\psi_{i}\}_{i=1}^{\infty}$ are linearly independent then $\langle\psi_{i},\psi\rangle_{\mathrm{L}^{2}(\mu)}\left(\beta-\sum_{k=1}^{q}\lambda_{k}\beta_{i}^{p_{k}}\right)=0$ for all $i\in\mathbb{N}$. Since $\psi\neq 0$, then at least one $\langle\psi_{i},\psi\rangle_{\mathrm{L}^{2}(\mu)}\neq 0$. For this $i$ we then must have $\beta=\sum_{k=1}^{q}\lambda_{k}\beta_{i}^{p_{k}}$ and so we are back in Case 1.
@@
In order to prove the proposition, we want to proceed as in [36, Theorem 1.2] where the authors show the analogous result for a single Laplacian matrix $\Delta_{n,\varepsilon_{n}}$. In particular, the proof relies on the following results:
@@
    the functional $\langle v,\Delta_{n,\varepsilon_{n}}v\rangle_{\mathrm{L}^{2}(\mu_{n})}$ $\Gamma$-converges to $\langle v,\Delta_{\rho}v\rangle_{\mathrm{L}^{2}(\mu)}$ [36, Theorem 1.4];
@@
    if a sequence satisfies $\sup_{n}\langle v,\Delta_{n,\varepsilon_{n}}v\rangle_{\mathrm{L}^{2}(\mu_{n})}\leq C$ and $\|v_{n}\|_{\mathrm{L}^{2}(\mu_{n})}\leq C$, then there exists a converging subsequence in $\mathrm{TL}^{2}(\Omega)$ [36, Theorem 1.4].
@@
    The fact that $\langle v,\mathcal{L}_{n}^{(q)}v\rangle_{\mathrm{L}^{2}(\mu_{n})}$ $\Gamma$-converges to $\langle v,\mathcal{L}^{(q)}v\rangle_{\mathrm{L}^{2}(\mu)}$ was shown in the proof of [85, Theorem 3.5].
@@
    If we assume that a sequence satisfies $\sup_{n}\langle v_{n},\mathcal{L}_{n}^{(q)}v_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}\leq C$ and $\|v_{n}\|_{\mathrm{L}^{2}(\mu_{n})}\leq C$, then in particular $\sup_{n}\langle v_{n},\Delta_{n,\varepsilon_{n}^{(q)}}^{p_{1}}v_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}\leq C$ and $\|v_{n}\|_{\mathrm{L}^{2}(\mu_{n})}\leq C$: we can therefore use [22, Theorem 2] to deduce the existence of a converging subsequence in $\mathrm{TL}^{2}(\Omega)$.
Specifically, let us start with the eigenvalues. First, we recall that since $\mathcal{L}_{n}^{(q)}$ is self-adjoint and positive semi-definite, we can apply the Courant-Fisher characterization of eigenvalues [16, Max-min theorem] to infer that
@@
Base case $i=1$
We first observe that the graphs $(\Omega_{n},W_{n,\varepsilon_{n}^{(k)}})$ are connected $\mathbb{P}$-a.e. for $n$ large enough. This follows from the ordering $\varepsilon_{n}^{(1)}>\cdots>\varepsilon_{n}^{(q)}$ and from Assumption (L.1), which guarantees connectivity in the random geometric graph regime [41, 64]. Consequently, for all sufficiently large $n$, the first eigenpair $(0,\mathbf{1})$, where $\mathbf{1}\in\mathbb{R}^{n}$ is the constant-one vector, is shared across all Laplacians $\Delta_{n,\varepsilon_{n}^{(k)}}$. Thus, the first eigenpair of the discrete operator $\mathcal{L}_{n}^{(q)}$ is given by $(\beta_{n,1},\psi_{n,1})=(0,\mathbf{1})$.
@@
Proof of the lower bound.   Let $S\in\Sigma_{i-1}$, where $\Sigma_{i-1}$ denotes the subspaces of $\mathrm{L}^{2}(\Omega)$ of dimension i1i-1. In this case, we will also write $S^{\perp}$ for the orthogonal complement of $S$ with respect to the inner product in $\mathrm{L}^{2}(\mu)$. Let $\{v_{1},\dots,v_{i-1}\}$ be an orthonormal basis of $S$. For each $\ell=1,\dots,i-1$, the $\limsup$-inequality in [85, Theorem 3.5] ensures the existence of a sequence of functions $v_{n,\ell}\in\mathrm{L}^{2}(\mu_{n})$ such that $(\mu_{n},v_{n,\ell})\to(\mu,v_{\ell})$ in $\mathrm{TL}^{2}(\Omega)$ as $n\to\infty$. By [36, Proposition 2.6], we have for all $1\leq\ell\leq i-1$,
@@
By (51) and [36, Proposition 2.6], it is straight-forward to check that $\tilde{v}_{n,\ell}\to v_{\ell}$ in $\mathrm{TL}^{2}(\Omega)$ for $1\leq\ell\leq i-1$. Let $S_{n}\in\Sigma_{n,i-1}$ be the subset spanned by $\{\tilde{v}_{n,1},\dots,\tilde{v}_{n,i-1}\}$.
@@
$\liminf_{n\to\infty}\min_{v\in S_{n}^{\perp},\,\|v|_{\mathrm{L}^{2}(\mu_{n})}=1}\langle\mathcal{L}_{n}^{(q)}v,v\rangle_{\mathrm{L}^{2}(\mu_{n})}<\infty$
and, without loss of generality (see [36, 87]), we can assume that
@@
By [22, Theorem 2], we therefore obtain a converging subsequence $(\mu_{n_{m}},w_{n_{m}})\to(\mu,w)$ in $\mathrm{TL}^{2}(\Omega)$. By [36, Proposition 2.6], we deduce that $\|w\|_{\mathrm{L}^{2}(\mu)}=\lim_{m\to\infty}\|w_{n_{m}}\|_{\mathrm{L}^{2}(\mu_{n_{m}})}=1$. Furthermore, since $w_{n_{m}}\in S_{n_{m}}^{\perp}$ and $\tilde{v}_{n_{m},\ell}\to v_{\ell}$, we also have $\langle w,v_{\ell}\rangle_{\mathrm{L}^{2}(\mu)}=\lim_{m\to\infty}\langle w_{n_{m}},\tilde{v}_{n_{m},\ell}\rangle_{\mathrm{L}^{2}(\mu_{n_{m}})}=0$ for $1\leq\ell\leq i-1$, which implies that $w\in S^{\perp}$. Combining the latter facts about $w$, we estimate as follows:
@@
where we used the $\liminf$-inequality of [85, Theorem 3.5] for (53) and (50) for (54). Finally, taking the supremum of all $S\in\Sigma_{i-1}$ in (52), applying the Courant-Fisher characterization to the self-adjoint and positive semi-definite operator $\mathcal{L}^{(q)}$ and using Lemma 4.6, we obtain

### v2

@@ 4.2.1 Convergence of truncated energies in the single Laplacian case
First, we tackle the $\liminf$-inequality. We assume that $(\nu,v)\in\mathrm{TL}^{2}(\Omega)$ and that $(\nu_{n},v_{n})\to(\nu,v)$ in $\mathrm{TL}^{2}$. If $\liminf_{n\to\infty}(\mathcal{S}\mathcal{J})_{n,\{\varepsilon_{n}\},K_{n}}^{(1,\{s\})}((\nu_{n},v_{n}))=+\infty$, then the inequality is trivial. Hence, without loss of generality, let us assume that $\sup_{n\in\mathbb{N}}(\mathcal{S}\mathcal{J})_{n,\{\varepsilon_{n}\},K_{n}}^{(1,\{s\})}((\nu_{n},v_{n}))\leq C$. In particular, this implies that $\nu_{n}=\mu_{n}$, $\langle\phi_{n,\varepsilon_{n},k},v_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}=0$ for all $k>K_{n}$ and $\sup_{n\in\mathbb{N}}\sum_{k=1}^{K_{n}}a_{n,\varepsilon_{n},k}^{s}\langle\phi_{n,\varepsilon_{n},k},v\rangle_{\mathrm{L}^{2}(\mu_{n})}^{2}\leq C$. Since we have $\mu_{n}\to\nu$ weakly (by the $\mathrm{TL}^{2}$-convergence assumption—see Proposition 2.2) and $\mu_{n}\to\mu$ weakly (convergence of the empirical measures), we conclude (by the uniqueness of weak limits) that $\nu=\mu$. We then proceed as in [Stuart, Theorem 2.2].
Let us start by assuming that $\sum_{k=1}^{\infty}\beta_{k}^{s}\langle v,\psi_{k}\rangle_{\mathrm{L}^{2}(\Omega)}^{2}<\infty$. In particular, since $\phi_{n,\varepsilon_{n},k}\to\psi_{k}$ and $v_{n}\to v$ in $\mathrm{TL}^{2}(\Omega)$ [Stuart], we have that $\langle\phi_{n,\varepsilon_{n},k},v_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}\to\langle v,\psi_{k}\rangle_{\mathrm{L}^{2}(\mu)}$ by [GARCIATRILLOS2018239, Proposition 2.6]. Furthermore, by [GARCIATRILLOS2018239, Theorem 1.2], we have $a_{n,\varepsilon_{n},k}\to\beta_{k}$. Now, let $\delta>0$ and pick $K$ such that
@@
For the $\limsup$-inequality, we let $(\nu,v)\in\mathrm{TL}^{2}(\Omega)$. If $(\mathcal{S}\mathcal{J})_{n,\{\varepsilon_{n}\},K_{n}}^{(1,\{s\})}((\nu,v))=\infty$, the inequality is trivial, so we assume that $\nu=\mu$ and $\sum_{k=1}^{\infty}\beta_{k}^{s}\langle\psi_{k},v\rangle_{\mathrm{L}^{2}(\mu)}^{2}<\infty$ or equivalently $v\in\mathrm{W}^{s,2}(\Omega)$ [Stuart]. If we can prove the $\limsup$-inequality on a dense subset of $\{\mu\}\times\mathrm{W}^{s,2}(\Omega)$, namely $\{\mu\}\times\mathrm{C}_{c}^{\infty}(\Omega)$, we can conclude due to [Trillos3, Remark 2.7].
@@
where we used the fact that the eigenvalues are ordered for (47) and [weihs2023consistency, Lemma 4.19] for (48). We know from [GARCIATRILLOS2018239, Theorem 1.4] that $T_{1}\to 0$ and from the proof of [Stuart, Theorem 2.2] that $a_{n,\varepsilon_{n},K_{n}+1}^{2}\to\infty$ which allows us to conclude that $(\mu_{n},\bar{v}_{n})\to(\mu,v)$ in $\mathrm{TL}^{2}(\Omega)$.
@@
where used [weihs2023consistency, Lemma 4.15] for (49), [weihs2023consistency, Proposition 4.21] and the fact that $T_{2}\to 0$ for (50). Squaring the last inequality, we obtain the $\limsup$-inequality.
@@
Let us now consider a sequence $(\mu_{n},v_{n})$ minimizing $(\mathcal{S}\mathcal{J})_{n,\{\varepsilon_{n}\},\Psi,K_{n}}^{(1,\{s\})}$ with $\sup_{n\in\mathbb{N}}\|v_{n}\|_{\mathrm{L}^{2}(\mu_{n})}\leq C$. In particular, we note that $\langle\phi_{n,\varepsilon_{n},k},v_{n}\rangle_{n}=0$ for all $k>K_{n}$ and recall that $K_{n}\to\infty$. Therefore, we can apply the same proof as in [Stuart, Theorem 2.2] to show that there exists a converging subsequence in $\mathrm{TL}^{2}(\Omega)$.
Specifically, $\sup_{n\in\mathbb{N}}\|v_{n}\|_{\mathrm{L}^{2}(\mu_{n})}\leq C$ implies that $\sup_{n\in\mathbb{N}}\sum_{k=1}^{K_{n}}\langle v_{n},\phi_{n,\varepsilon_{n},k}\rangle_{\mathrm{L}^{2}(\mu_{n})}^{2}\leq C.$ Hence, by a diagonal procedure, we can find a sequence $n_{m}\to\infty$ such that for every $k$, $\langle v_{n_{m}},\psi_{n_{m},\varepsilon_{n_{m}},k}\rangle_{\mathrm{L}^{2}(\mu_{n_{m}})}$ converges to some coefficient $\gamma_{k}$. By Fatou’s lemma, $\sum_{k=1}^{\infty}|\gamma_{k}|^{2}\leq\liminf_{m\to\infty}\sum_{k=1}^{n_{m}}|\langle v_{n_{m}},\phi_{n_{m},\varepsilon_{n_{m}},k}\rangle_{\mathrm{L}^{2}(\mu_{n_{m}})}|^{2}\leq C$, so we can define $v=\sum_{k=1}^{\infty}\gamma_{k}\psi_{k}\in\mathrm{L}^{2}(\mu)$. Using [Stuart, Lemma 7.7], we obtain a sequence $R_{n_{m}}\to\infty$ such that $\sum_{k=1}^{R_{n_{m}}}\langle v_{n_{m}},\phi_{n_{m},\varepsilon_{n_{m}},k}\rangle_{\mathrm{L}^{2}(\mu_{n_{m}})}\phi_{n_{m},\varepsilon_{n_{m}},k}\to v$ in $\mathrm{TL}^{2}(\Omega)$. We note that $R_{n_{m}}$ can always be picked such that $R_{n_{m}}\leq K_{n_{m}}$. Indeed, the $\mathrm{TL}^{2}$-convergence resulting from [Stuart, Lemma 7.7] holds for any sequence converging to $\infty$ and majorized by $R_{n_{m}}$: therefore, we can always pick $\tilde{R}_{n_{m}}=\min\{R_{n_{m}},K_{n_{m}}\}$.
@@
$\qquad\qquad+\|\sum_{k=1}^{R_{n_{m}}}\langle v_{n_{m}},\phi_{n_{m},\varepsilon_{n_{m}},k}\rangle_{\mathrm{L}^{2}(\mu_{n_{m}})}\phi_{n_{m},\varepsilon_{n_{m}},k}\circ T_{n_{m}}-v\|_{\mathrm{L}^{2}(\mu)}$
@@
$\qquad\qquad+\|\sum_{k=1}^{R_{n_{m}}}\langle v_{n_{m}},\phi_{n_{m},\varepsilon_{n_{m}},k}\rangle_{\mathrm{L}^{2}(\mu_{n_{m}})}\phi_{n_{m},\varepsilon_{n_{m}},k}\circ T_{n_{m}}-v\|_{\mathrm{L}^{2}(\mu)}$
@@
where the last inequality follows from the fact that $\sup_{n\in\mathbb{N}}\sum_{k=1}^{K_{n}}a_{n,\varepsilon_{n},k}^{s}\langle v_{n},\phi_{n,\varepsilon_{n},k}\rangle_{\mathrm{L}^{2}(\mu_{n})}^{2}\leq C$ since $(\mu_{n},v_{n})$ are minimizers of $(\mathcal{S}\mathcal{J})_{n,\{\varepsilon_{n}\},\Psi,K_{n}}$ (see also [weihs2023consistency, Lemma 4.25]). In order to conclude that $v_{n_{m}}$ to $v$ in $\mathrm{TL}^{2}(\Omega)$, we note the following two facts: the first term in the last inequality tends to 0 as argued in [Stuart, Theorem 2]; the second term tends to 0 since $\sum_{k=1}^{R_{n_{m}}}\langle v_{n_{m}},\psi_{n_{m},k}\rangle_{\mathrm{L}^{2}(\mu_{n_{m}})}\psi_{n_{m},k}\to v$ in $\mathrm{TL}^{2}(\Omega)$. By Proposition 2.6, we know that the limiting point $v$ is a minimizer of $(\mathcal{S}\mathcal{J})_{\infty,\Psi}^{(1,\{s\})}$. ∎
@@ 4.2.2 Spectral convergence of $\mathcal{L}_{n}^{(q)}$
$\mathcal{L}^{(q)}\psi_{i}=\sum_{k=1}^{q}\lambda_{k}\Delta_{\rho}^{p_{k}}\psi_{i}=\sum_{k=1}^{q}\lambda_{k}\beta_{i}^{p_{k}}\psi_{i}.$
@@
Again, as $\{\psi_{i}\}_{i=1}^{\infty}$ are linearly independent then $\langle\psi_{i},\psi\rangle_{\mathrm{L}^{2}(\mu)}\left(\beta-\sum_{k=1}^{q}\lambda_{k}\beta_{i}^{p_{k}}\right)=0$ for all $i\in\mathbb{N}$. Since $\psi\neq 0$, then at least one $\langle\psi_{i},\psi\rangle_{\mathrm{L}^{2}(\mu)}\neq 0$. For this $i$ we then must have $\beta=\sum_{k=1}^{q}\lambda_{k}\beta_{i}^{p_{k}}$ and so we are back in Case 1. ∎
@@
In order to prove the proposition, we want to proceed as in [GARCIATRILLOS2018239, Theorem 1.2] where the authors show the analogous result for a single Laplacian matrix $\Delta_{n,\varepsilon_{n}}$. In particular, the proof relies on the following results:
@@
    the functional $\langle v,\Delta_{n,\varepsilon_{n}}v\rangle_{\mathrm{L}^{2}(\mu_{n})}$ $\Gamma$-converges to $\langle v,\Delta_{\rho}v\rangle_{\mathrm{L}^{2}(\mu)}$ [GARCIATRILLOS2018239, Theorem 1.4];
@@
    if a sequence satisfies $\sup_{n}\langle v,\Delta_{n,\varepsilon_{n}}v\rangle_{\mathrm{L}^{2}(\mu_{n})}\leq C$ and $\|v_{n}\|_{\mathrm{L}^{2}(\mu_{n})}\leq C$, then there exists a converging subsequence in $\mathrm{TL}^{2}(\Omega)$ [GARCIATRILLOS2018239, Theorem 1.4].
@@
    The fact that $\langle v,\mathcal{L}_{n}^{(q)}v\rangle_{\mathrm{L}^{2}(\mu_{n})}$ $\Gamma$-converges to $\langle v,\mathcal{L}^{(q)}v\rangle_{\mathrm{L}^{2}(\mu)}$ was shown in the proof of [weihs2025Hypergraphs, Theorem 3.5].
@@
    If we assume that a sequence satisfies $\sup_{n}\langle v_{n},\mathcal{L}_{n}^{(q)}v_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}\leq C$ and $\|v_{n}\|_{\mathrm{L}^{2}(\mu_{n})}\leq C$, then in particular $\sup_{n}\langle v_{n},\Delta_{n,\varepsilon_{n}^{(q)}}^{p_{1}}v_{n}\rangle_{\mathrm{L}^{2}(\mu_{n})}\leq C$ and $\|v_{n}\|_{\mathrm{L}^{2}(\mu_{n})}\leq C$: we can therefore use [Stuart, Theorem 2] to deduce the existence of a converging subsequence in $\mathrm{TL}^{2}(\Omega)$.
Specifically, let us start with the eigenvalues. First, we recall that since $\mathcal{L}_{n}^{(q)}$ is self-adjoint and positive semi-definite, we can apply the Courant-Fisher characterization of eigenvalues [Chavel, Max-min theorem] to infer that
@@
Base case $i=1$.
We first observe that the graphs $(\Omega_{n},W_{n,\varepsilon_{n}^{(k)}})$ are connected $\mathbb{P}$-a.e. for $n$ large enough. This follows from the ordering $\varepsilon_{n}^{(1)}>\cdots>\varepsilon_{n}^{(q)}$ and from Assumption (L.1), which guarantees connectivity in the random geometric graph regime [goel, DBLP:books/ox/P2003]. Consequently, for all sufficiently large $n$, the first eigenpair $(0,\mathbf{1})$, where $\mathbf{1}\in\mathbb{R}^{n}$ is the constant-one vector, is shared across all Laplacians $\Delta_{n,\varepsilon_{n}^{(k)}}$. Thus, the first eigenpair of the discrete operator $\mathcal{L}_{n}^{(q)}$ is given by $(\beta_{n,1},\psi_{n,1})=(0,\mathbf{1})$.
@@
Proof of the lower bound.   Let $S\in\Sigma_{i-1}$, where $\Sigma_{i-1}$ denotes the subspaces of $\mathrm{L}^{2}(\Omega)$ of dimension i1i-1. In this case, we will also write $S^{\perp}$ for the orthogonal complement of $S$ with respect to the inner product in $\mathrm{L}^{2}(\mu)$. Let $\{v_{1},\dots,v_{i-1}\}$ be an orthonormal basis of $S$. For each $\ell=1,\dots,i-1$, the $\limsup$-inequality in [weihs2025Hypergraphs, Theorem 3.5] ensures the existence of a sequence of functions $v_{n,\ell}\in\mathrm{L}^{2}(\mu_{n})$ such that $(\mu_{n},v_{n,\ell})\to(\mu,v_{\ell})$ in $\mathrm{TL}^{2}(\Omega)$ as $n\to\infty$. By [GARCIATRILLOS2018239, Proposition 2.6], we have for all $1\leq\ell\leq i-1$,
@@
By (52) and [GARCIATRILLOS2018239, Proposition 2.6], it is straight-forward to check that $\tilde{v}_{n,\ell}\to v_{\ell}$ in $\mathrm{TL}^{2}(\Omega)$ for $1\leq\ell\leq i-1$. Let $S_{n}\in\Sigma_{n,i-1}$ be the subset spanned by $\{\tilde{v}_{n,1},\dots,\tilde{v}_{n,i-1}\}$.
@@
$\liminf_{n\to\infty}\min_{v\in S_{n}^{\perp},\,\|v\|_{\mathrm{L}^{2}(\mu_{n})}=1}\langle\mathcal{L}_{n}^{(q)}v,v\rangle_{\mathrm{L}^{2}(\mu_{n})}<\infty$
and, without loss of generality (see [GARCIATRILLOS2018239, weihs2023consistency]), we can assume that
@@
By [Stuart, Theorem 2], we therefore obtain a converging subsequence $(\mu_{n_{m}},w_{n_{m}})\to(\mu,w)$ in $\mathrm{TL}^{2}(\Omega)$. By [GARCIATRILLOS2018239, Proposition 2.6], we deduce that $\|w\|_{\mathrm{L}^{2}(\mu)}=\lim_{m\to\infty}\|w_{n_{m}}\|_{\mathrm{L}^{2}(\mu_{n_{m}})}=1$. Furthermore, since $w_{n_{m}}\in S_{n_{m}}^{\perp}$ and $\tilde{v}_{n_{m},\ell}\to v_{\ell}$, we also have $\langle w,v_{\ell}\rangle_{\mathrm{L}^{2}(\mu)}=\lim_{m\to\infty}\langle w_{n_{m}},\tilde{v}_{n_{m},\ell}\rangle_{\mathrm{L}^{2}(\mu_{n_{m}})}=0$ for $1\leq\ell\leq i-1$, which implies that $w\in S^{\perp}$. Combining the latter facts about $w$, we estimate as follows:
@@
where we used the $\liminf$-inequality of [weihs2025Hypergraphs, Theorem 3.5] for (54) and (51) for (55). Finally, taking the supremum of all $S\in\Sigma_{i-1}$ in (53), applying the Courant-Fisher characterization to the self-adjoint and positive semi-definite operator $\mathcal{L}^{(q)}$ and using Lemma 4.6, we obtain